Plank’s Plants had net income of $6,000 on sales of $70,000 last year. The firm paid a dividend of $600. Total assets were $300,000, of which $150,000 was financed by debt.

a.What is the firm’s sustainable growth rate?

b. If the firm grows at its sustainable growth rate, how much debt will be issued next year?

c. What would be the maximum possible growth rate if the firm did not issue any debt next year?
Answer #1

a) Payout ratio = Dividends / Net Income = 600 / 6000 = 10%

ROE = Net Income / Equity = 6,000 / 150,000 = 4.00%

Sustainable growth rate = ROE x (1 - payout ratio) = 4% x (1 - 10%) = 3.60%

b) New debt will grow by the same rate = 150,000 x 3.6% = $5,400

c) Internal Growth rate = ROA x b / (1 - ROA x b)

where, ROA = Net Income / Asset = 6000 / 300000 = 2%, b = Retention ratio = 1 - 10% = 90%

=> g = 2% x 90% / (1 - 2% x 90%) = 1.83%